find the equation of a line expressed in general form using a slope of -4/9 containing the points (-8,-10)
Answer by stanbon(75887)

find the equation of a line expressed in general form using a slope of -4/9 containing the points (-8,-10)
------------------------
Form: y = mx + b
-10 = (-4/9)(-8) + b
---------
-10 = (32/9) + b
-90 = 32 + 9b
9b = -122
b = -122/9
==================
Equation:
y = (-4/9)x - (122/9)
===========================
Cheers,
Stan H.